OpenStudy (callisto):

Coordination Geometry question #1

OpenStudy (callisto):

Perhaps this is easier, not sure :S

OpenStudy (anonymous):

Shouldnt the length between P and R be the same between S and Q?

OpenStudy (cwrw238):

r has coordinate (5+2), (-2+11) = (7,9)

OpenStudy (callisto):

@Romero Not really, it's just a //gram...

OpenStudy (cwrw238):

SLOPE OF pr = (7--8),9--4 = 15 / 13

OpenStudy (callisto):

@cwrw238 I got it!, thanks, though the slope is 13/15 as I got :P

OpenStudy (callisto):

But the coordinates for R is very useful! Good observation!

OpenStudy (anonymous):

How I saw it is that the shift from P to Q would be the same from S to R

OpenStudy (callisto):

slope = (9--4 ) / (7--8) = 13 / 15

OpenStudy (cwrw238):

OH YES - 13/15 - i got the subtraction wrong way round yw

OpenStudy (callisto):

I learnt a new trick :) Hooray ~~~

OpenStudy (cwrw238):

yea - coods of r is 2 to right of Q(x) and 11 up from Q(y)